OpenStudy (anonymous):

vertex and roots of f(t)-16t^2+70t+4

OpenStudy (anonymous):

f(t)=-16^2+70t+4 sorry

OpenStudy (dumbcow):

f'(t) should be -32t + 70 to find vertex set equal to 0

OpenStudy (anonymous):

ahhh sorry it's -16t^2+70+4 we're supposed to find -b/2a to find t, then plug that in the find vertex, and then the roots i came up with were 35+or-sqrt1097 divided by 16

OpenStudy (dumbcow):

sorry about that, vertex looks good check the sqare root part in the quadratic formula

OpenStudy (dumbcow):

i get 35 +- sqrt(1289) divided by 16

OpenStudy (anonymous):

b^2=35*35=1225 4ac=4(8)(4)=128 1225-128=1097? I think thats right, its asking for x-intercepts to graph it. how do i graph this?

OpenStudy (dumbcow):

redo the problem, this time looking closely at the signs.

OpenStudy (anonymous):

-4 got it. thank you!!!!!!!

OpenStudy (dumbcow):

4ac = 4(-8)(2)

OpenStudy (anonymous):

ooooh yeah. haha. thank you so much. but that means there are no x-intercepts right?

OpenStudy (dumbcow):

ok so now x = (35+- sqrt(1289))/16 those two x values are your x-intercepts graph them as 2 points. (x1,0) and (x2,0)

OpenStudy (dumbcow):

get a graphing calculator to check your answers

OpenStudy (anonymous):

but you can't graph that what we got for x

OpenStudy (dumbcow):

do you mean on a graphing calculator or by hand

OpenStudy (anonymous):

so like the question is about baseball and how long it takes for the ball to hit the ground and after how many seconds. when i graph it, (0,4) is the y-int, (35/16, 1289/16) is the vertex (or the highest point the ball gets to) and then to figure out when the ball lands on the ground would be the x-int

OpenStudy (dumbcow):

the x-int would be one of the answers derived from the quadratic formula.

OpenStudy (anonymous):

even if its like never ending full of decimals? would i just round it up?

OpenStudy (dumbcow):

yes round to two or three decimal places depending on what your teacher prefers.

OpenStudy (anonymous):

oh ok. i think i got it this time then. thanks for all your help! =]

OpenStudy (dumbcow):

you're welcome. next time you ask for help it would be beneficial to state the entire problem not just the part you're stuck on. it gives me context. Thanks!